# A coupled-channel perspective analysis on bottom-strange molecular pentaquarks

### v1

@@ II.1 The effective interactions
in which $\mathcal{M}(h_{1}h_{2}\to h_{3}h_{4})$ denotes the scattering amplitude for the $h_{1}h_{2}\to h_{3}h_{4}$ process and $m_{i}$ is the mass of the particle $h_{i}$. The Fourier transformation with respect to $\bm{q}$ leads to the effective potential in position space,
@@
Here, the parameter $\Lambda$ is introduced as an UV cut-off originates from the fact that the hadrons have a non-zero size to account for the inner structures of the interacting hadrons.
The corresponding one-boson-exchange effective potentialsare taken from Ref. Chen:2022svh and listed in Table 1, where $\mathcal{G}=-2$ for $I=1/2$ system and $\mathcal{G}=1$ for $I=3/2$ system. The explicit expressions for factors $\mathcal{F}_{1,2}$, $U$ and $Y$ in the effective potentials listed in Table 1 are given by
@@
where $S(\hat{r},\mathbf{a},\mathbf{b})\equiv 3(\hat{r}\cdot\mathbf{a})(\hat{r}\cdot\mathbf{b})-\mathbf{a}\cdot\mathbf{b}$. The values of relevant parameters are listed in Table 2 Liu:2011xc ; Chen:2017xat ; Kaymakcalan:1983qq .
@@
$l_{s}=2l_{B}$ 7.300
$g_{1}=(\sqrt{8}/3)g_{4}$ 1.000
$\beta_{s}g_{v}=-2\beta_{B}g_{v}$ 12.000
$\lambda_{s}g_{v}=-2\sqrt{2}\lambda_{I}g_{v}$ $19.200~{}\textrm{GeV}^{-1}$
$g_{\sigma}$ $-3.650$
$g$ 12.000
@@ II.2 Complex scaling method
At the present work, in order to obtain possible poles for these investigated systems, the complex scaling method (CSM) is applied Moiseyev:1998gjp ; Ho:1983lwa . In the CSM, the relative distance $\bm{r}$ and the conjugate momentum $\bm{p}$ are replaced by
$\bm{r^{\prime}}\to\bm{r}e^{i\theta},\bm{p^{\prime}}\to\bm{p}e^{-i\theta}$ (23)
where the scaling angle $\theta$ is chosen to bepositive. Applying such replacement to the Schrödinger equation, we get the complex scaled Schrödinger equation for the coupled channels which read
@@
It is worth noting that the properties of the solutions of the complex scaling Schrödinger equation can be predicted by the so-called ABC theorem Aguilar:1971ve ; Balslev:1971vb , which means
@@
    The wave functions for resonant states should be square-integrable on the complex plane, which is the same as bound state.
@@
According to this theorem, one can locate the poles on the complex plane. Moreover, in this work, the orbital wave functions are expanded in terms of a set of Gaussian basis functions. With the obtained wave functions, the root-mean-square (RMS) radii $r_{RMS}$ and component proportions $P$ can be calculated by T.yo ; Lin:2023ihj ; Shimizu:2016rrd
@@
It is worth to mention that the scaling angle $\theta$ should be larger than $1/2\text{Arg}(\Gamma/2E)$ to ensure the normalizability of wave functions of the resonant states Myo:2014ypa .
@@ III Results and discussions
Performing the above procedure, we can systematically investigate the bottom-strange molecular pentaquarks by solving coupled channel Schrödinger equation. In this work, the only free parameter is the UV cutoff $\Lambda$ in Eq. (21), which may vary for different coupled systems being investigated and it lays within the range of $800\sim 5000$ MeV. We firstly deal with bottom baryon $Y_{b}$ and $\bar{K}^{*}$ meson systems to reveal possible bound and resonant states and give them reasonable interpretations. The same technique can be applied in the analysis of bottom baryon $Y_{b}$ and $K$ meson systems. Our estimations for these investigated systems depending on the cutoff value $\Lambda$ are plotted in Figure 2 and listed in Table 3. In the present work, both $S-D$ wave mixing effects and coupled channel effects are taken into account. According to the isospin, spin, and parity, the bottom baryon and anti-strange meson systems can be classified as $1/2(1/2^{-})\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$, $3/2(1/2^{-})\Sigma_{b}\bar{K}/\Sigma_{b}\bar{K}^{*}$, $1/2(3/2^{-})\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$, and $3/2(3/2^{-})\Sigma_{b}\bar{K}^{*}$ channel, respectively. The corresponding classification also exists for the bottom baryon and strange meson systems.
@@
$I(J^{P})=1/2(3/2^{-})$ 1260 2.64 6511 (41.20 0.06 40.34 17.16 0.32 0.92)
1280 1.38 6506 (36.97 0.09 31.78 29.28 0.45 1.43
1300 0.98 6499 (33.72 0.11 25.67 38.26 0.51 1.73)
@@ III.1 Bottom baryon and anti-strange meson systems
In this subsection, we firstly discuss the coupled $I(J^{PC})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ systems. As shown in Fig. 2(a), one could find that a bound state and a resonance emerge within the range $\Lambda=880\sim 1100$ $\rm{MeV}$. When the cutoff $\Lambda$ is set to 880 $\rm{MeV}$, the bound state locates below $\Sigma_{b}\bar{K}$ threshold with binding energy about $4$ $\rm{MeV}$, the $r_{RMS}$ is 3 $\rm{fm}$ and dominated by the $\Sigma_{b}\bar{K}(^{2}S_{1/2})$ channel. Sliding the cutoff to $1080$ $\rm{MeV}$, the mass varies to be around 6222 $\rm{MeV}$ and the $r_{RMS}$ varies to be 0.8 $\rm{fm}$, which is consistent with the sizes of exotic hadronic molecular state. Thus, this bound state is a good candidate of the particle $\Xi_{b}(6227)$. These results favors the conclusion in Refs. Huang:2018bed ; Zhu:2020lza ; Mutuk:2024elj . Meanwhile, with the cutoff $\Lambda=970$ $\rm{MeV}$, we can obtain a resonant state with $E=6693-17i$ $\rm MeV$ and $r_{RMS}=1.70+1.43i$ $\rm{fm}$, which is mainly composed of the $\Sigma_{b}\bar{K}^{*}(^{2}S_{1/2})$ component. Also, it can be regarded as a good hadronic molecular state. Moreover, our predictions indicate that the pion exchange potential is crucial to form the resonance while the contribution from $\eta$ exchange interaction is negligible. For the $3/2(1/2^{-})$ $\Sigma_{b}\bar{K}^{*}$ state, when the cutoff $\Lambda$ lies in the range of $3360$ to $4100$ $\rm{MeV}$, a loosely bound state is found with $r_{\text{RMS}}$ varying between $4$ and $1$ $\rm{fm}$. However, this range for cutoff value is quite different from the empirical value of the deuteron, and then no molecular state is favored in the $3/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Sigma_{b}\bar{K}^{*}$ system.
Figure 3: The complex energy eigenvalues of $I(J^{P})=1/2(1/2^{-})$ system with varying the angle $\theta$ from $30^{\circ}\sim 40^{\circ}$.

### v2

@@ II.1 The effective interactions
in which $\mathcal{M}(h_{1}h_{2}\to h_{3}h_{4})$ denotes the scattering amplitude for the $h_{1}h_{2}\to h_{3}h_{4}$ process, and $m_{i}$ is the mass of the particle $h_{i}$. The Fourier transform with respect to $\bm{q}$ leads to the effective potential in position space,
@@
Here, the parameter $\Lambda$ is introduced as an UV cut-off originates from the fact that the hadrons have a non-zero size to account for the internal structure of the interacting hadrons.
The corresponding one-boson-exchange effective potentials are taken from Ref. Chen:2022svh and listed in Table 1, where $\mathcal{G}=-2$ for the $I=1/2$ system and $\mathcal{G}=1$ for the $I=3/2$ system. The explicit expressions for factors $\mathcal{F}_{1,2}$, $U$ and $Y$ in the effective potentials listed in Table 1 are given
@@
where $S(\hat{r},\mathbf{a},\mathbf{b})\equiv 3(\hat{r}\cdot\mathbf{a})(\hat{r}\cdot\mathbf{b})-\mathbf{a}\cdot\mathbf{b}$. The values of relevant parameters are listed in Table 2 Liu:2011xc ; Chen:2017xat ; Kaymakcalan:1983qq . The subscript labels $m_{\pi 1}$, $m_{\pi 2}$, etc., represent the effective masses of the exchanged mesons, which arise from the energy transfer $q^{0}$ during the Fourier transformation. Specifically, for the $Y(\Lambda_{2},m_{\pi 2},r)$-type potential, we use
$m_{\pi 2}$ $=$ $\sqrt{m_{\pi}^{2}-(q^{0})^{2}},\quad\Lambda_{2}=\sqrt{\Lambda^{2}-(q^{0})^{2}},$ (26)
$q^{0}$ $=$ $\frac{m_{1}^{2}+m_{4}^{2}-m_{2}^{2}-m_{3}^{2}}{2(m_{3}+m_{4})}=\frac{m_{\bar{K}}^{2}-m_{\bar{K}^{*}}^{2}}{2(m_{\Sigma_{b}}+m_{\bar{K}})}.$ (27)
For the $U(\Lambda_{1},m_{\pi 1},r)$-type potential, the definitions are
$m_{\pi 1}$ $=$ $\sqrt{(q^{0})^{2}-m_{\pi}^{2}},\quad\Lambda_{1}=\sqrt{\Lambda^{2}-(q^{0})^{2}},$ (28)
$q^{0}$ $=$ $\frac{m_{1}^{2}+m_{4}^{2}-m_{2}^{2}-m_{3}^{2}}{2(m_{3}+m_{4})}=\frac{m_{\Lambda_{b}}^{2}+m_{\bar{K}}^{2}-m_{\bar{K}^{*}}^{2}-m_{\Sigma_{b}}^{2}}{2(m_{\Sigma_{b}}+m_{\bar{K}})}.$ (29)
The same conventions apply to other exchanged mesons such as $\rho_{1,2}$, $\eta_{2}$, and $\omega_{1,2}$.
@@
$l_{s}=2l_{B}$ 7.3
$g_{1}=(\sqrt{8}/3)g_{4}$ 1.0
$\beta_{s}g_{v}=-2\beta_{B}g_{v}$ 12.0
$\lambda_{s}g_{v}=-2\sqrt{2}\lambda_{I}g_{v}$ $19.2~\textrm{GeV}^{-1}$
$g_{\sigma}$ 3.65-3.65
$g$ 12.00
@@ II.2 Complex scaling method
At present, in order to obtain possible poles for these investigated systems, the complex scaling method (CSM) is applied. In the CSM, the relative distance $\bm{r}$ and the conjugate momentum $\bm{p}$ are replaced by
$\bm{r^{\prime}}\to\bm{r}e^{i\theta},\quad\bm{p^{\prime}}\to\bm{p}e^{-i\theta}$ (30)
where the scaling angle $\theta$ is chosen to be positive. Applying such a replacement to the Schrödinger equation, we get the complex-scaled Schrödinger equation for the coupled channels which reads
@@
It is worth mentioning that the properties of the solutions of the complex scaling Schrödinger equation can be predicted by the so-called ABC theorem Aguilar:1971ve ; Balslev:1971vb , which means
@@
    The wave functions for resonant states should be square-integrable on the complex plane, which is the same as for a bound state.
@@
Figure 2: Schematic eigenvalue distributions of $H_{\theta}$ in the coupled-channel two-body systems.
According to this theorem, one pick out can the bound state or resonance from the continuum states as shown in Figure 2. This technique has been widely and successfully applied in quantum scattering theory and few-body systems. In our work, it provides a robust tool for identifying resonances in the coupled channel systems. Moreover, in this work, the orbital wave functions are expanded in terms of a set of Gaussian basis functions. With the obtained wave functions, the root-mean-square (RMS) radii $r_{RMS}$ and component proportions $P$ can be calculated by T.yo ; Lin:2023ihj
@@
Interestingly, the root mean square (RMS) radius $r_{\rm RMS}$ for the resonances can be a complex number. In such cases, one can use the interpretation scheme proposed by T. Berggren, which generalizes the concept of expectation values from bound states to resonances Berggren:1970wto . According to this scheme, the real part of the complex $r_{\rm RMS}$ represents the usual physical expectation value, while the imaginary part indicates a measure of uncertainty in the observation. Numerical calculations of $r^{2}$ have supported this generalized interpretation Gyarmati:1972yac ; 1997matrix . Besides, the $P$ involves the square of the complex-scaled wave function $\psi^{\theta}_{i}(\bm{r})$, rather than its modulus squared. Consequently, $P$ is not strictly positive definite. Depending on the nodal structure and the phase of the wave function, $P$ can take on positive or negative real values, or even have a complex part. Moreover, it is worth mentioning that the scaling angle $\theta$ should be larger than $1/2\text{Arg}(\Gamma/2E)$ to ensure the normalizability of the wave functions of the resonant states Myo:2014ypa .
@@ III Results and discussions
Figure 4: The complex energy eigenvalues of $I(J^{P})=1/2(1/2^{-})$ system by varying the angle $\theta$ from $30^{\circ}\sim 40^{\circ}$.
Figure 5: (Color online) The $r$ dependence of the deduced effective potentials with cutoff of 1000 $\rm{MeV}$ for the $S$-wave $\Sigma_{b}\bar{K}^{*}$ system with $I(J^{P})=1/2(1/2^{-})$.
Performing the above procedure, we can systematically investigate the bottom-strange molecular pentaquarks by solving the coupled channel Schrödinger equation. In this work, the only free parameter is the UV cutoff $\Lambda$ in Eq. (24), which may vary for different coupled systems being investigated, and it lies within the range of $800\sim 5000$ MeV Liu:2009qhy ; Chen:2019asm . A reasonable cutoff for deuteron studies is estimated to be around 1000 $\rm MeV$, but it can also vary among different scenarios Tornqvist:1993ng ; Tornqvist:1993vu . A loosely bound state at a cutoff around $1000$ $\rm MeV$ could be a promising molecular state candidate; thus, we summarize our predictions for bottom-strange pentaquark molecular state systems with cutoff $\Lambda$ in a range of $800\sim 1100$ $\rm MeV$ in Table 5.
We firstly deal with the bottom baryon $Y_{b}$ and $\bar{K}^{*}$ meson systems to reveal possible bound and resonant states and give them reasonable interpretations. The same technique can be applied in the analysis of the bottom baryon $Y_{b}$ and $K$ meson systems. Our estimations for these investigated systems depending on the cutoff value $\Lambda$ are plotted in Figure 3 and listed in Table 3 and 4. In the present work, both $S$-$D$ wave mixing effects and coupled channel effects are taken into account. According to the isospin, spin, and parity, the bottom baryon and anti-strange meson systems can be classified as $1/2(1/2^{-})\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$, $3/2(1/2^{-})\Sigma_{b}\bar{K}/\Sigma_{b}\bar{K}^{*}$, $1/2(3/2^{-})\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$, and $3/2(3/2^{-})\Sigma_{b}\bar{K}^{*}$ channels, respectively. The corresponding classification also exists for the bottom baryon and strange meson systems.
@@
$I(J^{P})=1/2(3/2^{-})$ 1320 2.02 6510 (70.16 0.09 0.37 27.76 0.32 1.30)
1340 0.96 6502 (52.39 0.11 0.54 44.70 0.43 1.83)
1360 0.70 6490 (43.51 0.12 0.58 53.39 0.45 1.95)
Table 4: The numerical results for the bottom baryon and anti-meson interactions with $I(J^{P})=1/2(1/2^{-})$ .
$\Lambda(\rm{MeV})$ $r_{RMS}(\rm{fm})$ $E(\rm{MeV})$ ($(\Sigma_{b}\bar{K}^{*}(^{2}S_{1/2})$ $\Sigma_{b}\bar{K}^{*}(^{4}D_{1/2})$)
$I(J^{P})=1/2(1/2^{-})$ 840 4.37 6703.34 (99.65 0.35)
970 1.16 6687.43 (99.15 0.85)
1100 0.78 6651.48 (98.83 1.17)
$\Lambda(\rm{MeV})$ $r_{RMS}(\rm{fm})$ $E(\rm{MeV})$ ($(\Sigma_{b}\bar{K}(^{2}S_{1/2})$ $\Sigma_{b}\bar{K}^{*}(^{2}S_{1/2})$ $\Sigma_{b}\bar{K}^{*}(^{4}D_{1/2})$)
$I(J^{P})=1/2(1/2^{-})$ 900 1.890.20i1.89-0.20i 6699.720.96i6699.72-0.96i (0.150.13i-0.15-0.13i $99.39+0.13i$ 0.760.00i0.76-0.00i)
1000 $1.12+0.00i$ 6680.000.07i6680.00-0.07i ($0.02+0.04i$ 99.080.05i99.08-0.05i $0.90+0.01i$)
1100 0.820.03i0.82-0.03i 6643.7812.99i6643.78-12.99i (1.13+0.29i-1.13+0.29i 100.640.62i100.64-0.62i $0.49+0.33i$)
$\Lambda(\rm{MeV})$ $r_{RMS}(\rm{fm})$ $E(\rm{MeV})$ $($$\Lambda_{b}\bar{K}^{*}(^{2}S_{1/2})$ $\Lambda_{b}\bar{K}^{*}(^{4}D_{1/2})$ $\Sigma_{b}\bar{K}^{*}(^{2}S_{1/2})$ $\Sigma_{b}\bar{K}^{*}(^{4}D_{1/2})$)
$I(J^{P})=1/2(1/2^{-})$ 920 $2.46+5.02i$ 6697.4833.62i6697.48-33.62i (16.57+2.89i-16.57+2.89i $0.28+1.26i$ 116.264.14i116.26-4.14i 0.030.01i0.03-0.01i)
$10$ 1.010.22i1.01-0.22i 6651.1042.66i6651.10-42.66i (31.38+10.64i-31.38+10.64i $0.95+1.67i$ 130.0313.15i130.03-13.15i $0.40+0.84i$)
1100 0.880.27i0.88-0.27i 6605.0331.83i6605.03-31.83i (13.67+28.50i-13.67+28.50i $1.46+0.43i$ 110.8229.83i110.82-29.83i $1.39+0.90i$)
$\Lambda(\rm{MeV})$ $r_{RMS}(\rm{fm})$ $E(\rm{MeV})$ ($\Sigma_{b}\bar{K}(^{2}S_{1/2})$ $\Lambda_{b}\bar{K}^{*}(^{2}S_{1/2})$ $\Lambda_{b}\bar{K}^{*}(^{4}D_{1/2})$ $\Sigma_{b}\bar{K}^{*}(^{2}S_{1/2})$ $\Sigma_{b}\bar{K}^{*}(^{4}D_{1/2})$)
$I(J^{P})=1/2(1/2^{-})$ 960 1.220.55i1.22-0.55i 6692.8417.29i6692.84-17.29i ($1.24+0.25i$ 2.994.37i-2.99-4.37i 1.15+0.68i-1.15+0.68i $102.41+3.33i$ $0.49+0.11i$)
1030 0.980.09i0.98-0.09i 6659.0518.99i6659.05-18.99i ($1.36+0.35i$ 5.99+0.17i-5.99+0.17i 0.46+1.34i-0.46+1.34i 104.282.40i104.28-2.40i $0.81+0.54i$)
1100 0.840.10i0.84-0.10i 6615.8015.50i6615.80-15.50i ($0.89+0.51i$ 4.14+7.39i-4.14+7.39i $0.40+1.11i$ 101.579.75i101.57-9.75i $1.28+0.74i$)
@@ III.1.1 $I(J^{P})=1/2(1/2^{-})$ with $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$
In this subsection, we firstly discuss the coupled $I(J^{PC})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ systems. As shown in Fig. 3(a), one could find that a bound state and a resonance emerge within the range $\Lambda=880\sim 1100$ $\rm{MeV}$. When the cutoff $\Lambda$ is set to 880 $\rm{MeV}$, the bound state is located below the $\Sigma_{b}\bar{K}$ threshold with a binding energy of about 44 $\rm{MeV}$, the $r_{RMS}$ is 3 $\rm{fm}$, and it is dominated by the $\Sigma_{b}\bar{K}(^{2}S_{1/2})$ channel. Sliding the cutoff to $1080$ $\rm{MeV}$, the mass becomes around 6222 $\rm{MeV}$ and the $r_{RMS}$ is 0.8 $\rm{fm}$, which is consistent with the sizes of exotic hadronic molecular states. Thus, this bound state is a good candidate for the particle $\Xi_{b}(6227)$. These results favor the conclusion in Refs. Huang:2018bed ; Zhu:2020lza ; Mutuk:2024elj .
Meanwhile, at a cutoff of $\Lambda=960~\rm{MeV}$, we obtain a quasi-bound state with complex energy $E=6693-17i~\rm{MeV}$ and root-mean-square radius $r_{\rm RMS}=1.22-0.55i~\rm{fm}$, which is predominantly composed of the $\Sigma_{b}\bar{K}^{*}(^{2}S_{1/2})$ channel. The binding of this state mainly arises from the strong interaction in the (closed) $\Sigma_{b}\bar{K}^{*}$ channel, while the nonzero width is induced via its coupling to lower open channels such as $\Lambda_{b}\bar{K}^{*}$. This structure is consistent with the nature of a Feshbach-type resonance and can be regarded as a good hadronic molecular candidate.
Besides, We illustrate the complex energy eigenvalues of the $I(J^{P})=1/2(1/2^{-})$ system while varying the angle $\theta$ from $30^{\circ}\sim 40^{\circ}$ in Figure 4. Specifically, the complex energy plane reveals the nature of poles: bound states lie on the real axis below threshold, while resonance appears as pole in the complex plane. The behavior of continuum states under varying scaling angle $\theta$ illustrates the separation between physical and unphysical states. Additionally, the clustering of poles near certain thresholds, such as $\Sigma_{b}\bar{K}$, $\Lambda_{b}\bar{K}^{*}$, and $\Sigma_{b}\bar{K}^{*}$ highlights the role of channel coupling and threshold effects. These patterns help identify the dominant configurations and offer insight into the underlying interaction dynamics.
Furthermore, our results highlight that the pion-exchange potential plays a crucial role in forming this resonance, whereas the contribution from the $\eta$-exchange interaction is negligible. As shown in Fig. 5, the $\eta$ exchange provides a much weaker attractive potential compared to the $\pi$ exchange. The dominant role of the pion exchange arises from its lighter mass and longer range, which enhance the attractive interaction between hadrons. In contrast, the $\eta$ meson is heavier, and its contribution is significantly suppressed. This physical picture is now clearly stated in the revised text. Moreover, according to our calculations, the resonance pole still appears even when the $\eta$-exchange potential is omitted, provided that the same cutoff is used.
III.1.2 $I(J^{P})=3/2(1/2^{-})$ with $\Sigma_{b}\bar{K}/\Sigma_{b}\bar{K}^{*}$
[…]
III.1.3 $I(J^{P})=1/2(3/2^{-})$ with $\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$
Besides, we investigate the $I(J^{P})=1/2(3/2^{-})$ $\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ channel for the $Y_{b}$ and $\bar{K}^{(*)}$ system, and the corresponding results are listed in Table 3. According to our estimations, at a cutoff of 1460 MeV, a bound state appears below the $\Lambda_{b}\bar{K}^{*}$ threshold. This state is mainly composed of the $\Sigma_{b}\bar{K}^{*}(^{4}S_{3/2})$ (75.6%), $\Sigma_{b}\bar{K}^{*}(^{4}D_{3/2})$ (7.5%), and $\Lambda_{b}\bar{K}^{*}(^{4}S_{3/2})$ (12%) components, and shows a strong dependence on the cutoff value. However, since this cutoff is inconsistent with the empirical value used for describing deuteron systems, and the corresponding $r_{\rm RMS}$ does not fall within the typical range for a molecular state, we conclude that the $I(J^{P})=1/2(3/2^{-})$ $\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ system is unlikely to form a hadronic molecular state.
III.1.4 $I(J^{P})=3/2(3/2^{-})$ with $\Sigma_{b}\bar{K}^{*}$